OpenStudy (anonymous):

Evaluate the limit: limx→∞ (4x+5)/3

geerky42 (geerky42):

What do you get?

OpenStudy (anonymous):

@geerky42 I honestly have no idea since there is no x on the bottom

geerky42 (geerky42):

What happened to the value of \(\dfrac{4x+5}{3}\) as x getting bigger and bigger?

OpenStudy (anonymous):

@geerky42 is it moving closer to 0?

geerky42 (geerky42):

really? What made you think that?

OpenStudy (anonymous):

hard to explain. Am I wrong? What's right? is it moving away from 0?

geerky42 (geerky42):

\(x→∞\) basically mean x getting closer and closer to infinity. So we have \(\displaystyle \lim_{x→∞}\dfrac{4x+5}{3}\), which mean the value of \(\dfrac{4x+5}{3}\) as x approaches to \(\infty\)

geerky42 (geerky42):

say x is start at 0. f(x) is 5/3, right? at x=1, f(x) = 9/3 = 3 at x=2, f(x) = 13/3 at x=3, f(x) = 17/3 Note how numerator grows?

OpenStudy (anonymous):

Correct go on

geerky42 (geerky42):

ok, what happen to f(x) when x is becoming infinity?

geerky42 (geerky42):

if we keeping going on: x=5, ... x=6, ... \(\vdots\)

OpenStudy (anonymous):

it becomes infinity?

geerky42 (geerky42):

Gotcha

geerky42 (geerky42):

So \(\displaystyle \lim_{x\rightarrow \infty}\dfrac{4x+5}{3} = \infty\)

OpenStudy (anonymous):

thank you very much!!!

geerky42 (geerky42):

welcome